# Periodic Cosmic String Formation and Dynamics

### v1

@@ 4.1 Local Vortex Confinement and Annihilation
Figure 16: A continuation to later times of the dynamics shown in Fig. 5 for the parameters $D=2$, $\rho_{i}/v=5$, $\sigma_{\rho}=10^{-2}\rho_{i}$, and $\mathcal{J}_{0}=0.1v^{2}\mu$. Left column, top row: Relevant normalized energy densities $\tilde{K}_{\rho}=K_{\rho}v^{-2}\mu^{-2}$, etc. through the instability saturation and vortex formation. Left column, center row: The total number of strings in the two-dimensional simulation domain of volume $(20/\mu)^{2}$. Left column, bottom row: The time difference $\tau_{i}$ between sequential instances of maximal number of strings in units of $\tau_{T}\equiv\pi/\dot{\vartheta}_{\text{sim}}^{\text{avg}}$, where $\dot{\vartheta}_{\text{sim}}^{\text{avg}}$ is the moving time-average (taken over an interval $T=8/\mu$) of the spatial average of the magnitude of the angular speed, $\langle|\dot{\theta}|\rangle$. Note that the offset of the ratio $\tau_{i}/\tau_{T}$ from unity is consistent with the uncertainties of our numerical methods. Right column: As for center row of the left column, but extending to much later times. Animation(s) of the dynamics for these parameters are available at ZenodoVideos .
@@
Before we close this subsection, it is beneficial to contrast the findings in the global $U(1)$ case presented here with those for the gauged $U(1)$ case studied in East:2022ppo ; East:2022rsi . In the gauged $U(1)$ case, the number of vortices grows almost monotonically after initial formation due to interactions between the vortices and the background field; by contrast, in the global $U(1)$ case here, the number of vortices actually oscillates periodically and slowly decreases over long timescales. This difference in behavior mainly comes from the fact that the global vortices are confined to the domain boundaries, and therefore annihilate much more efficiently as compared to the gauged case, whereas the gauged vortices can absorb almost all the energy from the background gauge field, and therefore persist. Particularly in the three-dimensional setting, the overall evolution of the string network after formation is different. In the gauged $U(1)$ case, the strings, via their interactions with the background gauge fields and each other, quickly cascade to a large number of short strings, whereas the strings in the global $U(1)$ case instead form and annihilate periodically.
@@ 4.2.1 Large Initial Displacement
The merger of counter-rotating regions is active on timescales $\gg\tau$, as can be seen in Fig. 18, lasting until a state consisting of one single connected region with $\dot{\theta}>0$ and another with $\dot{\theta}<0$ forms. This late-time merging may be expected because the domain boundaries carry energy densities proportional to their length (i.e., surface tension). [^27] Throughout this paragraph, wherever we refer to the “length” of a domain boundary or the “area” of a region, we are invoking the terminology applicable for two spatial dimensions merely for concreteness and clarity of presentation. These statements all apply equally to the case of three spatial dimensions, for which “area” and “volume”, respectively, would be the appropriate language. As the counter-rotating regions merge, the fully connected length of the merged domain boundaries will increase, while the total length of all domain boundaries in the box decreases. At the same time, the ratio of the areas of the co-rotating regions to the counter-rotating regions remains roughly constant and approximately unity. Moreover, throughout the domain-merger process the (time-averaged) total number of vortices in the domain decreases by only about a factor of two from early [$\mu t\sim\mathcal{O}(30)$] to late [$\mu t\sim\mathcal{O}(800)$] times; see Fig. 16. This slight reduction of vortex density is likely driven by the reduction of total length of the domain boundaries where vortices are born. Reducing this boundary size and decreasing the vortex density frees up energy, which is pumped into a scalar radiation field with characteristic wavenumber comparable to the mass of the radial mode: $k_{\rm{rad}}\sim\mu$. In Fig. 18, this manifests itself as perturbations in $\dot{\theta}$ with spatial scales of the order $\sim\mu^{-1}$, particularly visible in the rightmost two panels. We discuss this radiation in the next subsection in more detail.
We also advance a few plausible (but speculative) reasons for the disappearance at intermediate times of epochs where there are no vortices anywhere within the simulation volume, and the re-appearance of these epochs at very late times; see Fig. 16. As a preliminary statement, we note that the string formation and annihilation are initially both synchronized in the whole simulation volume. Then, at intermediate times, owing to spatial gradients in the $\theta$ and $\dot{\theta}$ fields across the simulation volume, phasing mismatches can develop that cause the formation times of individual vortices in different spatial locations to begin to move out of temporal synchronization with each other. In particular, at intermediate times, the lengths of the domain boundaries of various distinct counter-rotating regions may also become of unequal length, which may allow for unequal string lifetimes on different-length boundaries (i.e., a longer/larger boundary can mean it takes longer for a pair of opposite-vorticity strings to find each other to annihilate). However, once larger-scale order is returned to the fields due to mergers of the counter- and co-rotating regions (see Fig. 18), such phasing mismatches may relax, which could explain the reemergence of durations where very few strings are present in the simulation volume at late times (see the left column of Fig. 16). We have not, however, definitively established whether this rough qualitative mechanism is primarily responsible for the results we see in simulation. Note however that vortex production is always approximately locally periodic on the scale of individual co- and counter-rotating regions.
@@ 4.2.2 Small Initial Displacement
We show a state of the system with $\rho_{i}/v=\sqrt{2}+0.1$ in Fig. 19 that exhibits both clearly attached and potentially detached vortices. The attached vortex pairs (i.e., the pairs located in the $x<0$ part of the domain shown in Fig. 19) lie along a domain wall characterized by $\rho\ll v$ and $\kappa=-\partial_{\alpha}\theta\partial^{\alpha}\theta\gg\mu^{2}$. However, these domain boundaries no longer necessarily form closed loop-like structures in two-dimensions, suppressing efficient vortex–anti-vortex annihilation. Furthermore, this could allow vortices (formed earlier as pairs) to move off the domain boundaries and potentially behave (at least temporarily) similar to an isolated and non-relativistic vortex. In the $0<x\lesssim 5$ part of the domain shown in Fig. 19, two such possible examples can be seen. Although it is ambiguous what one means by “attached to” vs. “detached from” a domain boundary, [^28] Perhaps a more unambiguous approach would be to identify isolated points of non-zero vorticity that are separated from other points of non-zero vorticity by region that exhibits $\kappa=-\partial_{\alpha}\theta\partial^{\alpha}\theta<0$. Note however that this is a condition which the prospective candidates in Fig. 19 would fail. each member of this vortex–anti-vortex pair exhibits less evidence of being located on some long (length $\gg\mu^{-1}$) filamentary domain-boundary structure on which $\rho\ll v$; that is, at least as compared to the (anti-)vortices that live in the $x<0$ part of Fig. 19. Although the reduced initial field value assumed for the results shown in Fig. 19 may potentially allow vortices to detach from these more extended field structures, this does not necessarily mean that a larger number of vortices can survive in the simulation domain. Indeed, the total number of vortex pairs formed also scales down with the initial displacement (down to none, when $\rho_{i}=\rho_{c}$). In particular, as we detail next, vortex annihilation is still sufficiently efficient, even for $\rho_{i}/v=\sqrt{2}+0.1$, to completely annihilate all vortices. As such, any detachment effect that may exist is largely irrelevant for the overall dynamics of the system.
@@
In Fig. 20, we show the evolution of this same case from the onset of the parametric resonance through the end state in both two and three spatial dimensions. Let us focus first on the 2D results. Analogous to the large-initial-amplitude setting, vortex–anti-vortex pairs form on domain boundaries and annihilate on timescales $\tau$ associated with $\langle|\dot{\theta}|\rangle$ in the counter-rotating regions once the parametric resonance saturates. However, in stark contrast to the previously considered scenario, after a few $\tau$ this periodic behavior gives way to slow, but complete, annihilation of all vortices present in the simulation domain. In this case, the energy contained in vortices and boundaries of counter-rotating regions is efficiently injected into a scalar radiation state, such that no strings or domain boundaries remain after $\sim\mathcal{O}(100)\tau$. This occurs despite the partial detachment of some vortices from domain boundaries as described above. The final state reached here is one consisting purely of radiation. This manifests itself also in the distribution of energies shown in the top panel of Fig. 20: the $\tilde{K}_{\rho}\approx\tilde{K}_{\theta}\approx\tilde{G}_{\theta}>\tilde{G}_{\rho}$ and $\tilde{V}_{\rho}\ll 1$ behavior towards late times suggests that this radiation state is characterized by both massless $\theta$ modes and massive semi-relativistic $\rho$ modes.292929 The massless-ness of the $\theta$ modes can be seen from $\tilde{K}_{\theta}\approx\tilde{G}_{\theta}$ as would be expected from a dispersion relation $\omega_{\theta}^{2}=k_{\theta}^{2}$; the massive-ness of the $\rho$ modes can be seen from $\tilde{K}_{\rho}>\tilde{G}_{\rho}$, as would be expected from a dispersion relation $\omega_{\rho}^{2}=\mu^{2}+k_{\rho}^{2}$ for a semi-relativistic mode with mass $\mu$ and $k_{\rho}\lesssim\mu$. The results in 3D are analogous (except that strings form in loops on two-dimensional domain boundaries instead of vortex–anti-vortex pairs forming on one-dimensional domain boundaries).
@@ 4.3 String Network Endstate: Large Initial $\mathcal{J}_{0}$
We turn now to the case of large initial angular speeds, $\mathcal{J}_{0}\gg v^{2}\mu$, but with relatively small initial radial displacement: $\rho_{i}\gtrsim\rho_{c}$ (but also recall footnote 21). We focus on the system with $\mathcal{J}_{0}=10^{3}v^{2}\mu$, initial radial displacement $\rho_{i}/v=11.5$ and $\sigma_{\rho_{i}}=10^{-6}\rho_{i}$, in two spatial dimensions, $D=2$. For these parameters, we have $\rho_{c}\approx 11.3$ from (40).
@@ 5 Evolution in an Expanding Universe
along with its complex conjugate; here $H(t)\equiv\dot{a}/a$ is the Hubble parameter, $\nabla^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$ is the flat-space Laplacian in terms of comoving coordinates, and ˙$\dot{}\equiv\partial_{t}{}$. In the radiation-dominated setting, the Hubble parameter behaves as $H\propto a^{-(D+1)/2}$, implying a scale factor $a(t)\sim t^{2/(D+1)}$.
@@ 5.1 Preliminary Dependence on the Scale Factor
where, as before, $\mathcal{J}_{0}\equiv\rho_{0}^{2}\dot{\theta}_{0}$, which is covariantly conserved: i.e., $\mathcal{J}_{0}=(\text{const.})\times a^{-D}$. In addition to the $\rho_{i}$ initial conditions discussed above, we fix the rest of the initial conditions to be $\dot{\rho}_{0}(0)$, $\theta_{0}(0)=0$, and $\dot{\theta}_{0}(0)=\mathcal{J}_{0}/\rho_{i}^{2}$, depending on $\rho_{0}(0)=\rho_{i}$ (see discussion in footnote 33 below regarding the time $t=0$ in this context).
@@
During the initial growth phase of the parametrically unstable perturbations (i.e., while these modes remain sub-dominant), the primary coherent/background dynamics of the system can be understood using (44); see the detailed derivations of the following claimed scalings in App. D. In the regime of large initial radial amplitudes $\rho_{i}/v\gg 1$, but independently of $\mathcal{J}_{0}$ as long as $\sqrt{V_{\rm eff}^{\prime\prime}(\rho)}\gg H$ still holds, we have $\rho_{0}\propto a^{-D/3}$ initially and the relevant energy densities initially scale as $k_{\rho},k_{\theta},V_{\rho}\sim a^{-4D/3}$ for $D=2,3$. The case where $\mathcal{J}_{0}\gg v^{2}\mu$ and $\rho_{i}\approx\rho_{\text{min}}$ exhibits the same scalings. [^32] At $\mathcal{J}_{0}\gg\mu v^{2}$, we note that setting $\rho_{i}(a_{i})=\rho_{\text{min}}(a_{i})$ (i.e., the $\Phi$ field initial conditions are set such that the field instantaneously moves in a circular trajectory along the instantaneously circular degenerate potential minimum) does not necessarily result in having a non-oscillatory evolution of the $\rho$ field about the $a$-dependent minimum value $\rho_{\text{min}}\sim\mathcal{J}_{0}^{1/3}\propto a^{-D/3}$. This is because $\rho_{\text{min}}$ does not necessarily change adiabatically on the timescale of $1/\sqrt{V_{\rm eff}^{\prime\prime}(\rho_{\text{min}})}$, which in general prevents $\rho$ from tracking $\rho_{\text{min}}$ adiabatically. Instead, imposing our initial conditions even with $\rho_{i}(a_{i})=\rho_{\text{min}}(a_{i})$ will generally result in an oscillation of $\rho$ similar to that which would obtain if one took instead $\rho_{i}(a_{i})\neq\rho_{\text{min}}(a_{i})$, but still $\rho_{i}(a_{i})\sim\rho_{\text{min}}(a_{i})$. We do not comment further on this case. These scalings of the zero-mode would change were the radial-mode amplitude closer to the vacuum expectation value; that is $\rho_{i}\approx v$. In that case, and assuming $\mathcal{J}_{0}\lesssim v^{2}\mu$, we have $\rho_{0}-v\propto a^{-D/2}$ and the energy densities initially scale as $k_{\rho},V_{\rho}\sim a^{-D}$ and $k_{\theta}\sim a^{-2D}$.
@@
Once the parametric resonance drives $k$-modes into the nonlinear regime, the simplified picture of (44) no longer applies and the scalings of various energy components become determined by the late-time evolution of the system. Qualitatively, during this phase, the system reaches a quasi-stationary period when strings periodically form on the boundaries of the counter-rotating regions as they expand. During this phase, we observe that the field within the counter-rotating regions (CRR) rotates approximately circularly in field space, with constant velocity $\dot{\theta}_{\textsc{crr}}(a)$ at a fixed displaced $\rho_{\textsc{crr}}(a)$, with both quantities decreasing adiabatically as the scale factor $a$ increases. In the simulations, we also find that the total $j^{0}$ charge integrated over each counter-rotating region is approximately conserved, reflecting that the counter-rotating regions are of approximately constant comoving size, so that their physical volume grows as $a^{D}$. As the Universe expands, the $j^{0}$ charge density inside each CRR scales as [see also (3)]

### v2

@@ 4.1 Local Vortex Confinement and Annihilation
Figure 16: A continuation to later times of the dynamics shown in Fig. 5 for the parameters $D=2$, $\rho_{i}/v=5$, $\sigma_{\rho}=10^{-2}\rho_{i}$, and $\mathcal{J}_{0}=0.1v^{2}\mu$. Left column, top row: Relevant normalized energy densities $\tilde{K}_{\rho}=K_{\rho}v^{-2}\mu^{-2}$, etc. through the instability saturation and vortex formation. Left column, center row: The total number of strings in the two-dimensional simulation domain of volume $(20/\mu)^{2}$. Note that the overall scale for the total number of strings depends of course on the volume of the simulation domain; the important information presented here is how the number of strings in a fixed box changes in a relative sense over time. Left column, bottom row: The time difference $\tau_{i}$ between sequential instances of maximal number of strings in units of $\tau_{T}\equiv\pi/\dot{\vartheta}_{\text{sim}}^{\text{avg}}$, where $\dot{\vartheta}_{\text{sim}}^{\text{avg}}$ is the moving time-average (taken over an interval $T=8/\mu$) of the spatial average of the magnitude of the angular speed, $\langle|\dot{\theta}|\rangle$. Note that the offset of the ratio $\tau_{i}/\tau_{T}$ from unity is consistent with the uncertainties of our numerical methods. Right column: As for center row of the left column, but extending to much later times. Animation(s) of the dynamics for these parameters are available at ZenodoVideos .
@@
Before we close this subsection, it is beneficial to contrast the findings regarding the post-formation behavior of the strings in the global $U(1)$ case presented here with those for the gauged $U(1)$ case studied in East:2022ppo ; East:2022rsi . [^29] For the avoidance of doubt, we note that the string formation mechanisms in the global and gauged cases under discussion here are very different; we intend here to compare only the post-formation behavior of the strings. In the gauged $U(1)$ case, the number of vortices grows almost monotonically after initial formation due to interactions between the vortices and the background field; by contrast, in the global $U(1)$ case here, the number of vortices actually oscillates periodically and slowly decreases over long timescales. This difference in post-formation behavior mainly comes from the fact that the global vortices are confined to the domain boundaries, and therefore annihilate much more efficiently as compared to the gauged case, whereas the gauged vortices can absorb almost all the energy from the background gauge field, and therefore persist. Particularly in the three-dimensional setting, the overall evolution of the string network after formation is different. In the gauged $U(1)$ case, the strings, via their interactions with the background gauge fields and each other, quickly cascade to a large number of short strings, whereas the strings in the global $U(1)$ case instead form and annihilate periodically.
@@ 4.2.1 Large Initial Displacement
The merger of counter-rotating regions is active on timescales $\gg\tau$, as can be seen in Fig. 18, lasting until a state consisting of one single connected region with $\dot{\theta}>0$ and another with $\dot{\theta}<0$ forms. This late-time merging may be expected because the domain boundaries carry energy densities proportional to their length (i.e., surface tension). [^30] Throughout this paragraph, wherever we refer to the “length” of a domain boundary or the “area” of a region, we are invoking the terminology applicable for two spatial dimensions merely for concreteness and clarity of presentation. These statements all apply equally to the case of three spatial dimensions, for which “area” and “volume”, respectively, would be the appropriate language. As the counter-rotating regions merge, the fully connected length of the merged domain boundaries will increase, while the total length of all domain boundaries in the box decreases. At the same time, the ratio of the areas of the co-rotating regions to the counter-rotating regions remains roughly constant and approximately unity. Moreover, throughout the domain-merger process the (time-averaged) total number of vortices in the domain decreases by only about a factor of two from early $\mu t\sim\mathcal{O}(30)$] to late $\mu t\sim\mathcal{O}(800)$] times; see Fig. 16. This slight reduction of vortex density is likely driven by the reduction of total length of the domain boundaries where vortices are born. Reducing this boundary size and decreasing the vortex density frees up energy, which is pumped into a scalar radiation field with characteristic wavenumber comparable to the mass of the radial mode: $k_{\rm{rad}}\sim\mu$. In Fig. 18, this manifests itself as perturbations in $\dot{\theta}$ with spatial scales of the order $\sim\mu^{-1}$, particularly visible in the rightmost two panels. We discuss this radiation in the next subsection in more detail.
@@
We also advance a few plausible (but speculative) reasons for the disappearance at intermediate times of epochs where there are no vortices anywhere within the simulation volume, and the re-appearance of these epochs at very late times; see Fig. 16. As a preliminary statement, we note that the string formation and annihilation are initially both synchronized in the whole simulation volume. Then, at intermediate times, owing to spatial gradients in the $\theta$ and $\dot{\theta}$ fields across the simulation volume, phasing mismatches can develop that cause the formation times of individual vortices in different spatial locations to begin to move out of temporal synchronization with each other. In particular, at intermediate times, the domain boundaries of various distinct counter-rotating regions may also become of unequal length, which may allow for unequal string lifetimes on different-length boundaries (i.e., a longer/larger boundary can mean it takes longer for a pair of opposite-vorticity strings to find each other to annihilate). However, once larger-scale order is returned to the fields due to mergers of the counter- and co-rotating regions (see Fig. 18), such phasing mismatches may relax, which could explain the reemergence of durations where very few strings are present in the simulation volume at late times (see the left column of Fig. 16). We have not, however, definitively established whether this rough qualitative mechanism is primarily responsible for the results we see in simulation. Note however that vortex production is always approximately locally periodic on the scale of individual co- and counter-rotating regions.
@@ 4.2.2 Small Initial Displacement
We show a state of the system with $\rho_{i}/v=\sqrt{2}+0.1$ in Fig. 19 that exhibits both clearly attached and potentially detached vortices. The attached vortex pairs (i.e., the pairs located in the $x<0$ part of the domain shown in Fig. 19) lie along a domain wall characterized by $\rho\ll v$ and $\kappa=-\partial_{\alpha}\theta\partial^{\alpha}\theta\gg\mu^{2}$. However, these domain boundaries no longer necessarily form closed loop-like structures in two-dimensions, suppressing efficient vortex–anti-vortex annihilation. Furthermore, this could allow vortices (formed earlier as pairs) to move off the domain boundaries and potentially behave (at least temporarily) similar to an isolated and non-relativistic vortex. In the $0<x\mu\lesssim 5$ part of the domain shown in Fig. 19, two such possible examples can be seen. Although it is ambiguous what one means by “attached to” vs. “detached from” a domain boundary, [^31] Perhaps a more unambiguous approach would be to identify isolated points of non-zero vorticity that are separated from other points of non-zero vorticity by region that exhibits $\kappa=-\partial_{\alpha}\theta\partial^{\alpha}\theta<0$. Note however that this is a condition which the prospective candidates in Fig. 19 would fail. each member of this vortex–anti-vortex pair exhibits less evidence of being located on some long (length $\gg\mu^{-1}$) filamentary domain-boundary structure on which $\rho\ll v$; that is, at least as compared to the (anti-)vortices that live in the $x<0$ part of Fig. 19. Although the reduced initial field value assumed for the results shown in Fig. 19 may potentially allow vortices to detach from these more extended field structures, this does not necessarily mean that a larger number of vortices can survive in the simulation domain. Indeed, the total number of vortex pairs formed also scales down with the initial displacement (down to none, when $\rho_{i}=\rho_{c}$). In particular, as we detail next, vortex annihilation is still sufficiently efficient, even for $\rho_{i}/v=\sqrt{2}+0.1$, to completely annihilate all vortices. As such, any detachment effect that may exist is largely irrelevant for the overall dynamics of the system.
@@
In Fig. 20, we show the evolution of this same case from the onset of the parametric resonance through the end state in both two and three spatial dimensions. Let us focus first on the 2D results. Analogous to the large-initial-amplitude setting, vortex–anti-vortex pairs form on domain boundaries and annihilate on timescales $\tau$ associated with $\langle|\dot{\theta}|\rangle$ in the counter-rotating regions once the parametric resonance saturates. However, in stark contrast to the previously considered scenario, after a few $\tau$ this periodic behavior gives way to slow, but complete, annihilation of all vortices present in the simulation domain. In this case, the energy contained in vortices and boundaries of counter-rotating regions is efficiently injected into a scalar radiation state, such that no strings or domain boundaries remain after $\sim\mathcal{O}(100)\tau$. This occurs despite the partial detachment of some vortices from domain boundaries as described above. The final state reached here is one consisting purely of radiation. This manifests itself also in the distribution of energies shown in the top panel of Fig. 20: the $\tilde{K}_{\rho}\approx\tilde{K}_{\theta}\approx\tilde{G}_{\theta}>\tilde{G}_{\rho}$ and $\tilde{V}_{\rho}\ll 1$ behavior towards late times suggests that this radiation state is characterized by both massless $\theta$ modes and massive semi-relativistic $\rho$ modes.323232 The massless-ness of the $\theta$ modes can be seen from $\tilde{K}_{\theta}\approx\tilde{G}_{\theta}$ as would be expected from a dispersion relation $\omega_{\theta}^{2}=k_{\theta}^{2}$; the massive-ness of the $\rho$ modes can be seen from $\tilde{K}_{\rho}>\tilde{G}_{\rho}$, as would be expected from a dispersion relation $\omega_{\rho}^{2}=\mu^{2}+k_{\rho}^{2}$ for a semi-relativistic mode with mass $\mu$ and $k_{\rho}\lesssim\mu$. The results in 3D are analogous (except that strings form in loops on two-dimensional domain boundaries instead of vortex–anti-vortex pairs forming on one-dimensional domain boundaries).
@@ 4.3 String Network Endstate: Large Initial $\mathcal{J}_{0}$
We turn now to the case of large initial angular speeds, $\mathcal{J}_{0}\gg v^{2}\mu$, but with relatively small initial radial displacement: $\rho_{i}\gtrsim\rho_{c}$ (but also recall footnote 23). We focus on the system with $\mathcal{J}_{0}=10^{3}v^{2}\mu$, initial radial displacement $\rho_{i}/v=11.5$ and $\sigma_{\rho_{i}}=10^{-6}\rho_{i}$, in two spatial dimensions, $D=2$. For these parameters, we have $\rho_{c}\approx 11.3$ from (40).
@@ 5 Evolution in an Expanding Universe
along with its complex conjugate; here $H(t)\equiv\dot{a}/a$ is the Hubble parameter, $\nabla^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$ is the flat-space Laplacian in terms of comoving coordinates, and $\dot{}\equiv\partial_{t}{}$. In the radiation-dominated setting, the Hubble parameter behaves as $H\propto a^{-(D+1)/2}$, implying a scale factor $a(t)\sim t^{2/(D+1)}$.
@@ 5.1 Preliminary Dependence on the Scale Factor
where, as before, $\mathcal{J}_{0}\equiv\rho_{0}^{2}\dot{\theta}_{0}$, which is covariantly conserved: i.e., $\mathcal{J}_{0}=(\text{const.})\times a^{-D}$. In addition to the $\rho_{i}$ initial conditions discussed above, we fix the rest of the initial conditions to be $\dot{\rho}_{0}(0)$, $\theta_{0}(0)=0$, and $\dot{\theta}_{0}(0)=\mathcal{J}_{0}/\rho_{i}^{2}$, depending on $\rho_{0}(0)=\rho_{i}$ (see discussion in footnote 36 below regarding the time $t=0$ in this context).
@@
During the initial growth phase of the parametrically unstable perturbations (i.e., while these modes remain sub-dominant), the primary coherent/background dynamics of the system can be understood using (44); see the detailed derivations of the following claimed scalings in App. D Turner:1983he . In the regime of large initial radial amplitudes $\rho_{i}/v\gg 1$, but independently of $\mathcal{J}_{0}$ as long as $\sqrt{V_{\rm eff}^{\prime\prime}(\rho)}\gg H$ still holds, we have $\rho_{0}\propto a^{-D/3}$ initially and the relevant energy densities initially scale as $k_{\rho},k_{\theta},V_{\rho}\sim a^{-4D/3}$ for $D=2,3$. The case where $\mathcal{J}_{0}\gg v^{2}\mu$ and $\rho_{i}\approx\rho_{\text{min}}$ exhibits the same scalings. [^35] At $\mathcal{J}_{0}\gg\mu v^{2}$, we note that setting $\rho_{i}(a_{i})=\rho_{\text{min}}(a_{i})$ (i.e., the $\Phi$ field initial conditions are set such that the field instantaneously moves in a circular trajectory along the instantaneously circular degenerate potential minimum) does not necessarily result in having a non-oscillatory evolution of the $\rho$ field about the $a$-dependent minimum value $\rho_{\text{min}}\sim\mathcal{J}_{0}^{1/3}\propto a^{-D/3}$. This is because $\rho_{\text{min}}$ does not necessarily change adiabatically on the timescale of $1/\sqrt{V_{\rm eff}^{\prime\prime}(\rho_{\text{min}})}$, which in general prevents $\rho$ from tracking $\rho_{\text{min}}$ adiabatically. Instead, imposing our initial conditions even with $\rho_{i}(a_{i})=\rho_{\text{min}}(a_{i})$ will generally result in an oscillation of $\rho$ similar to that which would obtain if one took instead $\rho_{i}(a_{i})\neq\rho_{\text{min}}(a_{i})$, but still $\rho_{i}(a_{i})\sim\rho_{\text{min}}(a_{i})$. We do not comment further on this case. These scalings of the zero-mode would change were the radial-mode amplitude closer to the vacuum expectation value; that is $\rho_{i}\approx v$. In that case, and assuming $\mathcal{J}_{0}\lesssim v^{2}\mu$, we have $\rho_{0}-v\propto a^{-D/2}$ and the energy densities initially scale as $k_{\rho},V_{\rho}\sim a^{-D}$ and $k_{\theta}\sim a^{-2D}$.
@@
Once the parametric resonance drives $k$-modes into the nonlinear regime, the simplified picture of (44) no longer applies and the scalings of various energy components become determined by the late-time evolution of the system. Qualitatively, during this phase, the system reaches a quasi-stationary period when strings periodically form on the boundaries of the counter-rotating regions as they expand. During this phase, we observe that the field within the counter-rotating regions rotates approximately circularly in field space, with constant velocity $\dot{\theta}_{\textsc{crr}}(a)$ at a fixed displaced $\rho_{\textsc{crr}}(a)$, with both quantities decreasing adiabatically as the scale factor $a$ increases. In the simulations, we also find that the total $j^{0}$ charge integrated over each counter-rotating region is approximately conserved, reflecting that the counter-rotating regions are of approximately constant comoving size, so that their physical volume grows as $a^{D}$. As the Universe expands, the $j^{0}$ charge density inside each CRR scales as [see also (3)]